### v3

# OntoPortal-Astro, a Semantic Artefact Catalogue for Astronomy

Abstract
The astronomy communities are widely recognised as mature communities for their open science practices. However, while their data ecosystems are rather advanced and permit efficient data interoperability, there are still gaps between these ecosystems. Semantic artefacts (SAs) —e.g., ontologies, thesauri, vocabularies or metadata schemas— are a means to bridge that gap as they allow to semantically described the data and map the underlying concepts. The increasing use of SAs in astronomy presents challenges in description, selection, evaluation, trust, and mappings. The landscape remains fragmented, with SAs scattered across various registries in diverse formats and structures — not yet fully developed or encoded with rich semantic web standards like OWL or SKOS — and often with overlapping scopes. Enhancing data semantic interoperability requires common platforms to catalog, align, and facilitate the sharing of FAIR (Findable, Accessible, Interoperable and Reusable) SAs. In the frame of the FAIR-IMPACT project, we prototyped a SA catalogue for astronomy, heliophysics and planetary sciences. This exercise resulted in improved vocabulary and ontology management in the communities, and is now paving the way for better interdisciplinary data discovery and reuse. This article presents current practices in our discipline, reviews candidate SAs for such a catalogue, presents driving use cases and the perspective of a real production service for the astronomy community based on the OntoPortal technology, that will be called OntoPortal-Astro.
keywords:
Astronomy , Heliophysics , Planetary Sciences , Semantic artefact , Vocabularies , Ontologies , Thesaurus , Terminologies , Open Science , Semantic Artefact Catalogue , OntoPortal , Semantic Web , Metadata
journal: Astronomy &\&& Computing
[…]
1 Introduction
The set of resources in the so-called “virtual observatory” (Genova et al., 2000; Hill et al., 2009; Hanisch et al., 2015; Erard et al., 2018, 2020; Mampaey et al., 2025) forms a network of related data products, catalogues, services and tools allowing researchers to discover, access and process data for their research. One example is the International Virtual Observatory Alliance [^1] https://ivoa.net (IVOA) framework (Arviset et al., 2018), which proposes a registry (Demleitner et al., 2019) of resources gathering endpoints for services, datasets and catalogues, and is used by tools to discover and provide access to resources. This astronomy framework is an example of an open science ecosystem.
1.1 Astronomy communities
The astronomy community is mainly composed of three semantics sub-communities: celestial astronomy [^2] In this paper, we use the term “celestial astronomy” to refer to any topics outside the Solar System. Note that it’s not a commonly accepted definition for this concept, but we use it for the sake of simplicity. (objects are referenced to with their fixed sky coordinates, e.g., stars, galaxies, cosmology, etc); heliophysics (the study of the Sun, and the plasma environments throughout the Solar System); planetary sciences (the study of the Solar System objects, e.g., planets, comets, asteroids, etc). Each of these sub-communities have developed interoperability frameworks and semantic ecosystems, which are rather siloed up to now, compromising the semantic and technical interoperability of those ecosystems.
The main astronomy sub-community is the celestial astronomy community. Its data management infrastructure is organised around the IVOA, which is interoperability driven. The IVOA was founded in 2002, and is now composed of 23 national groups (at the time of writing), which act as relays between the community developments and the international cooperation level. The IVOA gathers twice per year with the so-called “IVOA Interop” conferences. With this fully bottom-up framework, the IVOA defines schemas, protocols and vocabularies, which are respectively managed by the Data Models, Data Access Layers, and Semantics working groups (WGs). In addition, a registry (managed by the IVOA Registry WG) provides a catalogue of IVOA compliant resources. Hence, this enables a stable and well-established ecosystem adopted by the vast majority of astronomy archives, data centres and tool developers. As a result, scientists are using applications, such as Aladin (Bonnarel et al., 2000) or TOPCAT (Taylor, 2005, 2011), which are fully relying on IVOA tooling without seeing any detail of the underlying interoperability layers.
The second sub-community is the heliophysics community. It is organised around the International Heliophysics Data Environment Alliance [^3] https://ihdea.net (IHDEA, Masson et al., 2024), which is distribution [^4] In the sense of DCAT (Data Catalog Vocabulary), see https://www.w3.org/TR/vocab-dcat-3/#Class:Distribution driven. The IHDEA was formed in 2018 by nine institutions (from laboratories to space agencies) willing to foster collaboration and standardisation of the heliophysics data ecosystem. However, the most of heliophysics community standards predates the inception of the IHDEA. The Space Physics Archive Search and Extract (SPASE, Roberts et al., 2018) is the main heliophysics information model, which development started in 2002. It is maintained by the SPASE Metadata Working Team. It defines a rather complete and evolutive way to describe space physics observational datasets, model runs, instruments or repositories. For each dataset, it is possible to define the various ways of the accessing the data. It is used by several major heliophysics data centres (NASA/SPDF, CNES/CDPP, IUGONET, see e.g., Abe et al., 2014; Clarke, 2016; Roberts et al., 2018) for their internal data model. Other metadata guidelines are available to this community (such as the ISTP metadata for data stored in CDF (Whipple and Lancaster, 1995; Candey, 1998), or the SOLARNET metadata for data stored in FITS (Hanisch et al., 2001; Haugan and Fredvik, 2021)). Data access protocols are also standardised, with, e.g., the HAPI protocol (Weigel et al., 2021). The IHDEA ecosystem is not as interoperable as the IVOA one. While some tools are providing a very simple access to heliophysics datasets (such as the AMDA [^5] https://amda.irap.omp.eu (Génot et al., 2021) or the CDAWeb [^6] https://cdaweb.gsfc.nasa.gov web interfaces), using IHDEA services or resources often require a deep understanding of interoperability interfaces and data formats. However ongoing developments are aiming at building tighter links with IVOA and improve open science in heliophysics.
The third sub-community is the planetary science community. It is mainly organised around the International Planetary Data Alliance [^7] https://planetarydata.org (IPDA), which is archive driven. The IPDA was formed in 2006 and is now composed of 13 members, which are mostly space agencies. The primary role of the IPDA is to develop and maintain the NASA Planetary Data System (PDS) (Hughes et al., 2017), which is an advanced information model for planetary exploration science archives. The PDS is used by all major space agency archive with planetary exploration missions: NASA/PDS, ESA/PSA, JAXA, ISRO, etc. Moreover, planetary scientists studying the planetary surfaces are using the OGC888 https://www.ogc.org/ (Open Geospatial Consortium) ecosystem, developed by the geoscience community (Earth observation). In 2021, a small group of planetary scientist formed the Planetary Domain Working Group, which role is to adopt the OGC standards for the study of planetary surfaces. Technical interoperability between OGC and IVOA has been studied for some years (Hare et al., 2018; Marmo et al., 2018; Minin et al., 2019; Erard and the VESPA team, 2024). However, semantic interoperability still needs to be implemented.
Each astronomy sub-community data framework implements various pieces of an open science ecosystem with a different goal: celestial astronomy focuses on interoperability; heliophysics, on distribution; and planetary sciences, on long term archive. This also implies that different ways of registering, connecting, publishing and accessing resources and their metadata.
1.2 Linked-data architecture
The linked-data architecture has been defined at the inception times of the World Wide Web (Berners-Lee and Hendler, 2001). The original framework — called RDF: Resource Description Framework — (Schreiber and Raimond, 2014) developed and published by the W3C proposes a description of resources and their relations, based on triples: a subject, a predicate and an object. The triples can be seen as sentences, the predicate being the verb. The set of triples creates an oriented and tagged graph of nodes. Nodes are resources (referred to as URI: Unified Resource Identifier). Predicates (or properties) designate the links connecting the ressources. The classes, their instances and the properties are usually defined in standard vocabularies or ontologies capturing a certain level of semantics that a machine will then rely to query and reason over the data.
In the present case, this graph of nodes, metadata and relations is capturing knowledge about scientific resources, hence we call it a “scientific knowledge graph”. Keeping this graph of resources usable and interoperable while it grows, requires semantic and technical coordination and synchronisation, so that terms are globally understood by the tools and actors providing the content and enabling the access. This semantic and technical coordination is achieved by methods, technologies and tools coming from a domain called, the Semantic Web for which semantics artefacts, are at the core (Corcho et al., 2024). A semantic artefact [^9] In Corcho et al. (2024), a “semantic artefact” is defined as a machine-actionable formalisation (represented using appropriate formats and serialisations, including RDF and non-RDF standards) of a conceptualisation, enabling sharing and reuse by humans and machines. In this paper, the term is used to refer to both machine-actionable and non-machine-actionable resources. (hereafter referred to as SA) — a broader term to include ontologies, terminologies, taxonomies, thesauri, vocabularies, metadata schemas — can be a simple list of terms (properties, classes of resources), or a more complex ontology allowing inference and logics between terms. We know that a unique absolute ontology is utopia. Each SAs encode contextual knowledge, which need to be understandable, shareable and reusable by humans and machines.
Without fully implementing yet the standard semantic web toolkit, the astronomy communities are implementing open linked-data frameworks. They all define resource classifications, propose community registries that provide links between the resources, and associate these with rich metadata. Upgrading these open science frameworks with semantic web technologies is then a matter of defining and mapping concepts without breaking the current solutions.
1.3 Semantics Artefacts for FAIR Astronomy
In Europe, the development of open data frameworks has been supported by the European Commission framework programs for many years. A significant number of astronomy projects have been funded throughout the last two decades. Table 1 presents an already long (though non-exhaustive) list of such grants, ranked by starting years. The ESCAPE project, which ended in 2023, has initiated the merging of astronomy and particle physics infrastructures and standards.
Acronym Name Topics Identifier Programme Time range
AVO Astrophysical virtual observatory A HPRI-CT-2001-50030 FP5 2001-2004
EUROPLANET European Planetology Network PS 1637 FP6 2005-2008
EUROVO-DCA The European virtual observatory data centre alliance A 31675 FP6 2006-2008
EuroVO-AIDA Euro-VO Astronomical Infrastructure for Data Access A 212104 FP7 2008-2010
EUROPLANET-RI European Planetology Network Research Infrastructure H,PS 228319 FP7 2009-2012
HELIO The Heliophysical Integrated Observatory H 238969 FP7 2009-2012
VAMDC Virtual Atomic and Molecular Data Center A,H,PS 239108 FP7 2009-2012
CASSIS Coordination Action for the integration of Solar System Infrastructures and Science H,PS 261618 FP7 2010-2013
EuroVO-ICE Euro-VO International Cooperation Empowerment A 261541 FP7 2010-2012
IMPEx Integrated Medium for Planetary Exploration H 262863 FP7 2011-2015
ESPAS Near-Earth Space Data Infrastructure for e-Science 283676 FP7 2011-2015
ASTERICS Astronomy ESFRI and Research Infrastructure Cluster A doi:10.3030/653477 H2020 2015-2019
EPN2020-RI Europlanet 2020 Research Infrastructure H,PS doi:10.3030/654208 H2020 2015-2019
ESCAPE European Science Cluster of Astronomy & Particle physics ESFRI research infrastructures A,H,$P$ doi:10.3030/824064 H2020 (EOSC) 2019-2023
SOLARNET Integrating High Resolution Solar Physics H doi:10.3030/824135 H2020 2019-2023
EPN-2024-RI Europlanet 2024 Research Infrastructure H,PS doi:10.3030/871149 H2020 2020-2024
PITHIA-NRF Plasmasphere Ionosphere Thermosphere Integrated Research Environment and Access services: a Network of Research Facilities H doi:10.3030/101007599 H2020 2021-2025
FAIR-IMPACT Implementing FAIR data and services in the European Open Science Cloud A,H,M,PS doi:10.3030/101057344 HORIZON (EOSC) 2022-2025
OSTRAILS Open Science Plan-Track-Assess Pathways A,H,M,PS,$P$ doi:10.3030/101130187 HORIZON (EOSC) 2024-2017
Table 1: European commission grants including astronomy data management developments and research infrastructures. The topics are listed with the following code: A - Astronomy; H - Heliophysics; M - Multi-disciplinary (outside astronomy); $P$ - Particle Physics; PS - Planetary Sciences. The identifiers before the H2020 programme are just links to the CORDIS web page for the project.
The scope of this study covers the astronomy community along with its three sub-domains (celestial astronomy, planetary sciences and heliophysics). Recent publications (e.g., Hurlburt and Timmons, 2022; Berriman, 2024; Cecconi, 2024; Landais et al., 2024; Servillat et al., 2024; Masson et al., 2024) show that each of these communities are claiming that they are producing a FAIR (Findable, Accessible, Interoperable and Reusable, Wilkinson et al., 2016) ecosystem. The evaluation of the FAIRness of astronomy digital resources is discussed and studied by many groups, and is out of the scope of this paper. However, one of the FAIR criteria [^10] I2: “(meta)data use vocabularies that follow FAIR principles” explicitly requires that community vocabularies also complies with the FAIR principles. Moreover, the fact that each sub-community is still siloed is a strong hint that things can be improved.
In 2018, the IVOA Solar System interest group started to bring heliophysics and planetary sciences needs in the IVOA ecosystem. The uptake of IVOA protocols in heliophysics and planetary sciences is growing, but at lot remains to be done, in particular on the semantic interoperability side, so that it is more widely adopted. This means that we need to harmonise our SAs, so that communities can exchange data and transform these data into knowledge — semantically described, interoperable, actionable, and open. Moreover, the adoption of common SAs across communities will improve the FAIRness of data products, since the metadata would be described with the same terms, or with terms that are related one to the other. The increasing use of SAs in astronomy presents challenges including their description, selection, evaluation, trustworthiness and mapping. The landscape remains fragmented, with SAs scattered across various registries in diverse formats and structures — not yet fully developed or encoded with rich semantic web standards based on RDF like OWL (Ontology for the Web Language, Bechhofer et al. (2004)) or SKOS (Simple Knowledge Organisation System, Miles and Bechhofer (2009)) — and often with overlapping scopes. There is thus need for a common place to catalogue, manage, assess, select and align FAIR SAs from several communities (Gray, 2015; Ritschel et al., 2016; Rovetto, 2023).
2 Current semantic artefact management in astronomy
The celestial astronomy community data alliance is the IVOA, which maintains an operational interoperability framework used by data repositories and science application platforms throughout the world. In this community, SAs are composed of vocabularies and schemas. The Semantics WG of the IVOA is managing the vocabularies used in the IVOA standards, following the Vocabularies in the VO specification (Demleitner et al., 2023). The vocabularies are listed in a dedicated web page [^11] https://ivoa.net/rdf/ where they are accessible using IVOA as well as with semantic web standard formats. Updates of the vocabularies are managed through a Vocabulary Enhancement Proposal (VEP) process (for details, see Demleitner et al., 2023). The maintenance of the vocabularies is managed in a dedicated Github repository [^12] https://github.com/ivoa-std/Vocabularies. The Data Model WG of the IVOA is managing the schemas, following the Virtual Observatory Data Model Language specification (Lemson et al., 2018), but they are not available in semantic web compliant formats. The role on SAs in the IVOA is to ensure the interoperability, with controlled vocabularies to be used in protocol and metadata fields. In the recent years, most of the controlled valued lists, which were previously maintained in schemas by the Data Model WG, have been transferred to the Semantics WG. In this way, the IVOA SAs have already been prepared for a wider use.
In the planetary science community, two main frameworks co-exist with different scopes. The IPDA maintains an advanced data archiving metadata schema for planetary exploration datasets, the Planetary Data System version 4 information model131313 https://pds.nasa.gov/datastandards/ (PDS4-IM) (Hughes et al., 2017). The foundations of the PDS4-IM rely on the OAIS Standard (Open Archive Information System, ISO 14721:2012). It is developed using semantic web related tools, such as Protégé (Musen, 2015), but is not implemented using RDF. The PDS4-IM is maintained and curated by a Change Control Board [^14] https://pds-engineering.jpl.nasa.gov/pds4/ccb/, which is supported from a Data Design WG [^15] https://pds-engineering.jpl.nasa.gov/pds4/ddwg/ representing the various PDS4 disciplinary nodes. The Planetary Domain WG of the OGC (Open Geospatial Consortium) proposes a set of standards and tools used for the study of planetary surfaces. Although the OGC already implements SAs in the form of RDF SAs (see, e.g., the ESIP —Earth Science Information Partners— Community Ontology Repository [^16] http://cor.esipfed.org), the Planetary Domain WG is not providing its SAs in RDF yet.
The heliophysics data ecosystem is managed by the IHDEA. This community is proposing a set of tools and standards for finding and accessing datasets for the heliophysics community. Semantic artefacts in this community were historically of three kinds: the SPASE Ontology (available as an XML schema), which includes lists of terms, properties and classes for defining various objects and resources (Persons, Observatories, Instruments, Datasets, Repositories, etc); the VSO (Virtual Solar Observatory) metadata; and the SOLARNET set of keywords (dedicated to Solar observations). Each of these three metadata standard are community effort to improve the findability and accessibility of their research products and their distributions. However, they are still currently not fully connected, and none of the them share an RDF implementation.
The IVOA, IPDA and IHDEA alliances are all worldwide collaborative groups, consensus and bottom-up driven, and based on best effort contributions. Interdisciplinary links between these communities have been mostly developed thanks to the Europlanet/VESPA [^17] http://www.europlanet-vespa.eu infrastructure (Erard et al., 2018). They focus on discoverability and implementation of plugins to extend the capabilities of existing tools. The semantic interoperability across the sub-communities approach started only recently, with the ongoing development of three common SAs: a vocabulary for “observation facilities” (Cecconi et al., 2023) and a proposal to the IHDEA to adopt two IVOA SAs for managing reference frames (Weigel et al., 2025a) and units (Weigel et al., 2025b).
In this paper, we present the results of a prototyping exercise. Thanks to the FAIR-IMPACT project, we have implemented an instance of the OntoPortal technology (Jonquet et al., 2023) to deploy a prototype SA catalogue for astronomy. Hereafter, we present current practices in astronomy, and review candidate SAs for such a cataloguing exercise. We report the use cases that have been driving this development. We present the first results, discuss the impact and outcomes, and layout a roadmap for the deployment of a real production service for the astronomy community based on the OntoPortal technology, that will be called OntoPortal-Astro. Throughout the paper, the astronomy OntoPortal instance and the OntoPortal-Astro terms refer respectively to the prototype instance development within the FAIR-IMPACT project, and the upcoming official server.
Among the vocabularies that we included in the astronomy OntoPortal instance, we can list the following items: the Unified Astronomy Thesaurus [^18] https://astrothesaurus.org/ (Frey and Accomazzi, 2018) (UAT); the IVOA vocabularies (including properties and terms for the IVOA protocols interoperability); the EPNcore (Europlanet-VESPA core metdata) dictionary (Erard et al., 2022); the IHDEA community metadata including the SPASE schema [^19] https://spase-group.org/data/, the Virtual Solar Observatory data model [^20] http://docs.virtualsolar.org/wiki/DataModel18; the IUGONET schema [^21] http://www.iugonet.org/product/metadata.jsp (Inter-university Upper atmosphere Global Observation NETwork); the SOLARNET metadata (Haugan and Fredvik, 2021); the CEF (Cluster Exchange Format) metadata (Allen et al., 2009); the IPDA information model [^22] https://pds.nasa.gov/datastandards/documents/im/. Other SAs will be included, such as: the VAMDC (Virtual Atomic and Molecular Data Centre) metadata schema [^23] http://dictionary.vamdc.eu/, the OGC planetary frame registry [^24] http://voparis-vespa-crs.obspm.fr. The GEMET (GEneral Multilingual Environmental Thesaurus) thesaurus [^25] https://www.eionet.europa.eu/gemet/ could also be included, however we might federate with to EarthPortal (Earth sciences community OntoPortal instance) for this SA. Regarding particle physics, the CERN Open Data [^26] https://github.com/cernopendata terms and schemas will also be included.
The two main challenges of this project are: (i) to identify the SAs, which are currently scattered in many places, and (ii) to produce SAs in a standard semantic web form (typically OWL, or SKOS). Besides being a requirement for being ingested into the astronomy OntoPortal instance, it is also a logical evolution for our artefacts. Indeed, most of the current SAs (except those from the celestial astronomy community) are in diverse forms, from lists of terms embedded in XML schemas, to unformatted lists of metadata in specification documents (PDF files, HTML pages, etc.) This “semantic lifting” shall be done by the semantics-related working groups or authorities of the relevant communities (e.g., the IVOA Semantics WG or the IHDEA dedicated teams), with the support of expert in semantics and partners from the OntoPortal Alliance (mostly from the AgroPortal team, the Agri-food community OntoPortal instance developed by INRAE) as well as from the OntoPortal-Astro team (see Section 6).
3 Use Cases
A series of use cases have been identified to trigger the developments of this SA catalogue dedicated to the astronomy communities.
3.1 Exoplanets
Exoplanetary objects (defined as planets orbiting around other stars than the Sun) and their environments are studied with increasing resolution and capabilities. Comparisons with solar system object observations become very promising. Solar system observations should then be available and semantically compliant with the IVOA astronomy ecosystem. Currently, the IVOA object type vocabulary (objtyp, see Table 1 for all namespace definitions) lists the term objtyp:planet defined as an “Extra-solar Planet" (not including the solar system objects). On the other hand, the EPNcore dictionary includes a list of target classes (tarcls), where the terms tarcls:planet and tarcls:exoplanet respectively refer to “solar system planets” and “exoplanets”. Similarly the UAT (uat) proposes two concepts: uat:1260 and uat:498 for “Solar system planet” and “Exoplanets” respectively. The IVOA UAT vocabulary (ivoauat) mirrors the UAT terms that thus also proposes the terms ivoauat:solar-system-planet and ivoauat:exoplanets. This requires careful management of the metadata across domains, with adequate mappings.
Currently, the uat and ivoauat thesauri are published as SKOS vocabularies. The objtyp and tarcls vocabularies are draft SAs prepared respectively by the IVOA Semantics WG, and by the Solar System interest group of the IVOA with the Europlanet team, and are being developed as OWL ontologies, using either Protégé, or the IVOA vocabulary framework. Mapping these SAs is not a difficult task, but collecting them into a single OntoPortal instance makes it easier.
The current IVOA use of vocabularies is strongly implicit: the IVOA specifications list which SA should be used in controlled-valued fields of protocols and interfaces. The IVOA client software are built with this implicit knowledge, which is enabling a technical interoperability. The semantic mapping between terms, if required for a use case, have thus to be implemented in clients, upstream from the IVOA interfaces and protocol queries. The mappings are currently stored in the IVOA vocabularies: e.g., in the IVOA UAT, each term is semantically mapped to its twin concept in the UAT thesaurus, using a skos:exactMatch property.
3.2 Space Weather
A widely accepted definition of Space Weather is: “Space Weather is the physical and phenomenological state of natural space environments. The associated discipline aims, through observation, monitoring, analysis and modelling, at understanding and predicting the state of the Sun, the interplanetary and planetary environments, and the solar and non-solar driven perturbations that affect them; and also at forecasting and nowcasting the possible impacts on biological and technological systems(Lilensten and Belehaki, 2009). Space weather is thus about heliophysics and its impact to the Earth environment. Space Weather services are being developed by “classical” weather organisations, such as the Met Office (UK), using mainly Earth observation software infrastructure (OGC), whereas the original data are produced by the heliophysics and astronomy communities (IHDEA and IVOA ecosystems). Collating and mapping SAs between OGC and the astronomy data ecosystems is required to efficiently exchange information, including mappings and coordination on the development of common ontologies. This implies mapping terms about reference frames, units of measure, physical quantities, sensor descriptions, etc.
3.3 Observation context
Modern instruments are getting very versatile with many modes of operations, thanks to the increasing digital processing capabilities, closer to the data acquisition devices, and implying automated data processing. Provenance metadata are increasingly implemented to trace these processing steps, requiring ontologies to efficiently model observational and instrumental contexts. Astronomy, heliophysics or particle physics ontologies need to be aligned and used in a coordinated manner to allow reuse in interdisciplinary studies.
The W3C has adopted a data model (Prov-DM) (Belhajjame et al., 2013) and an ontology (Prov-O) (Lebo et al., 2013) for managing provenance metadata. The IVOA inspired from those recommendations with the VO-PROV data model (Servillat et al., 2020), extending the W3C Prov-DM. The IPDA and IHDEA have not yet standardised such metadata. Mapping VO-PROV to Prov-DM (and thus Prov-O) is straightforward, but should be implemented. We should also explore how the IPDA and IHDEA teams are individually managing their provenance metadata when applicable. Note that for the general domain-agnostic semantic web vocabularies, the OntoPortal FAIR-IMPACT partners are currently working on a specific portal (LovPortal) to host the content of the Linked Open Vocabulary platform.
3.4 Harmonisation of nomenclature
Some concepts are used by all sub-communities of astronomy, such as the observed physical quantities, the observation facilities, the instrument types, or the coordinate reference frames. Harmonisation of these generic terms and concepts is needed for semantic interoperability. Such developments already started for observed parameters (Cecconi et al., 2014), observation facilities (Cecconi et al., 2023), as well as coordinate reference frames in an IHDEA working group. For observation facilities and reference frames, the current plan is to let the IVOA Semantics WG manage the merged vocabularies. The astronomy community OntoPortal instance will provide a platform to follow up and further develop these projects. Within OntoPortal every SA is automatically mapped one-another (at the lexical level) and term reuses are explicitly exposed and detected by the tool.
3.5 Scientific Knowledge Graphs
Connecting research outputs to persons, institutions and funders is a key task for enabling the FAIR principles, discovering research products or measuring the scientific impact of the projects and infrastructures. Several tools are implementing so called SKG, which can mean either Scholarly Knowledge Graph (such as with the OpenAIRE Graph infrastructure [^27] https://graph.openaire.eu), or Scientific Knowledge Graph (such as with the SciLake project [^28] https://scilake.eu). Moreover, the astronomy community registries, such as the IVOA registry or the SPASE registry, can also be named SKGs as they relate all the elements of the research ecosystem (data products, repositories, scholarly articles, institutions, persons, etc.)
Building efficient and FAIR knowledge graphs requires to use well-defined and cross-domain ontologies, so that term URIs can be used as search keys in queries (see, e.g., the “exoplanet” use case). OntoPortal-Astro will implements a very rich metadata model to describe SAs — based on the MOD metadata vocabulary [^29] https://github.com/FAIR-IMPACT/MOD (Dutta et al., 2017) — which will support the formal and semantic description of astronomy SAs in SKGs. Plus, it will embed a FAIRness assessment tool based on the O’FAIRe methodology (Ontology FAIRness Evaluation, Amdouni et al., 2022) to help in the selection of SAs. OntoPortal-Astro will thus facilitate the reuse of astronomy SAs making them more findable.
3.6 Models and observations
Astronomy is an observation-based science, since the environmental conditions of the observed objects and phenomena can’t be controlled. It is thus necessary to try and model the possible environmental conditions leading to the observations. Within the IVOA, the celestial astronomy community has developed the Simulation Data Model (Lemson et al., 2012), which can describe the astrophysical simulation parameters. However, due to its complexity, it has not been adopted by many teams. In the heliophysics community, the need to coupling measurements to simulations has been identified as a key issue (Masson et al., 2024), either to discover and reuse observational data as input for environment models, or to compare observations to output parameters of those models. Hence, the observed and modelled physical parameters must be described the same way. The same reasoning applies to planetary environment, such as the global atmospheric models of planets, or the planetary magnetosphere models (see, e.g., Erard et al., 2018; Achilleos et al., 2019; Azria et al., 2021; Trompet et al., 2022).
Describing observable properties (or scientific variables) within ontologies or other types of SAs, is a growing practice. It is present in multiple OntoPortal community instances such as Earth sciences (EarthPortal), Agri-food (AgroPortal) or Ecology (EcoPortal), using the Extensible Observation Ontology (OBOE, Madin et al., 2007) or the InteroperAble Descriptions of Observable Property Terminology [^30] https://i-adopt.github.io (I-ADOPT)). For instance, in EcoPortal, the ENVTHES (Environmental Thesaurus) already use the I-ADOPT representation to encode their scientific variable within these SAs  [^31] See, e.g., http://ecoportal.lifewatch.eu/ontologies/ENVTHES/30209.
4 First Results
Based on the use-cases presented in the previous section, work has been initiated mostly in the frame of the FAIR-IMPACT project.
4.1 Review of pre-exisiting Semantic Artefact Catalogues
A recent landscape analysis of catalogues containing SAs has identified three among the astronomy community: the IVOA vocabularies, the NASA/PDS data standards and the ESPAS Vocabulary broker (Jonquet and Grau, 2024b). SA catalogues have been classified according to their typology and the level of service offered. A listing corresponds to a simple online listing of SAs, while a library represents structured online listing and the a repository is considered as an advanced web application (Jonquet and Grau, 2024a). The IVOA vocabulary [^32] https://ivoa.net/rdf/ web page is considered as a SA library, which gathers the vocabularies developed and maintained by the IVOA Semantics WG. The NASA/PDS data standard [^33] https://pds.nasa.gov/datastandards web page is a SA listing, which provides all version of the NASA/PDS information model, and the associated vocabularies. The ESPAS Vocabulary Broker, which is consider as retired (Jonquet and Grau, 2024a), and is not accessible anymore at the time of writing, is a SA repository. It was providing access to the terms from the SPASE schema, the ESPAS project ontology, the GEMET ontology and the UAT thesaurus. Moreover, the ESPAS vocabulary used to offer mapping between terms of ontologies among the repository. Among those, solely the IVOA vocabulary library partly meets the SA catalogues FAIRness criteria proposed in Jonquet and Grau (2024a). Hence, no pre-existing solution satisfies the needs for a common astronomy SA catalogue. There is thus a need of harmonising and synchronising the different initiatives to avoid technical and semantic silos.
4.2 OntoPortal-Astro, an astronomy community OntoPortal
In the FAIR-IMPACT projet, an astronomy SA catalogue use-case has been developed (Cecconi et al., 2024). An OntoPortal instance for astronomy has been installed:
https://ontoportal-astro.eu [^34] At the time of writing, this server presents the astronomy OntoPortal prototype developed during the FAIR-IMPACT project, but will be replaced by the OntoPortal-Astro within a few months.
The goal of the prototype was two-fold: testing the technology to expose, and assess SAs in astronomy. We selected about 40 vocabularies and ontologies that we would use as tests of the usage of the framework, see Table 2. In addition to all the IVOA Vocabularies, we also drafted OWL ontology versions of several community SAs (three schemas and one vocabulary): (i) the NASA/PDS4 information model, (ii) the IHDEA/SPASE data model, (iii) the VAMDC schema, and (iv) the CERN Open Data terms. The development of an ontology version of a schema is more than a format conversion, since it implies a full modelling of the concepts, classes and properties, as well as minting URIs for the terms. Hence, the development of these SAs have been used to exercise the astronomy team members. We interacted with the schema maintainers to share our explorations. The NASA/PDS4, IHDEA/SPASE and VAMDC teams expressed a strong interest in our approach. In the future, we envision the sharing in OntoPortal-Astro as part of the life cycle of each of these SAs as it is advocate by classic vocabulary development methodologies such as Linked Open Terms (LOT, Poveda-Villalón et al., 2022) that has also include FAIR assessment steps to ensure the developed artefacts are FAIR by design and appropriately shared.
Name Metrics Topics Status Notes
Astronomical Subject Keywords Concepts: 394 A,H,PS Test Old ontology developed by the AAS (American Astronomical Society), still used by some journals.
CDPP TREPS tool Reference frames Classes: 78 H,PS Develop Intermediate ontology extracted from TREPS tool source code. To be merged into IVOA RefFrame
EPNcore Dataproduct Types Classes: 14 H,PS Develop EPNcore v2.0 Dataproduct Type vocabulary.
EPNcore Small Body Dynamical Classes Classes: 19 PS Develop EPNcore v2.0 Small Body Dynamical Class vocabulary.
EPNcore Spatial Frame Types Classes: 7 H,PS Develop EPNcore v2.0 Spatial Frame Type vocabulary.
EPNcore Target Classes Classes: 13 H,PS Develop EPNcore v2.0 Target Class vocabulary.
GEMET Thesaurus Concepts: 5739 H,PS Test Test import of the GEMET Thesaurus. To be coordinated with EarthPortal.
IVOA Content types of VO resources Classes: 22 M Ready Generic types of resources in the Virtual Observatory.
IVOA Datalink core Properties: 22 M Ready Properties linking described object to other resources, in the DataLink IVOA standard.
IVOA Object Types Classes: 151 A Develop Astronomical object types used in the IVOA, still in development.
IVOA Messengers Classes: 9 A,H,PS Ready Messenger (type of particule) providing the information observed in astronomy.
IVOA Product Types Concepts: 20 A Ready Data Product type vocabulary for the IVOA.
IVOA Reference Positions Classes: 6 A Ready Reference positions used in the IVOA.
IVOA Relation Types Properties: 14 M Ready Used for relations between resources in the IVOA registry
IVOA Roles of dates Properties: 14 M Ready Roles of dates in the IVOA Registry
IVOA Semantics Properties: 4 M Ready This vocabulary defines several properties used in the management of IVOA vocabularies.
IVOA Time Scales Classes: 9 A,H,PS Ready Time scales used in the IVOA.
IVOA Unified Astronomy Thesaurus Concepts: 2373 A,H,PS Ready IVOA Rendering of Unified Astronomy Thesaurus (UAT)
IVOA Unified Content Descriptors (UCD) v1.5 Classes: 6 Individuals: 554 M,A,H,PS Develop Rendering of IVOA UCD vocabulary as an ontology.
NASA NAIF Spacecraft Codes Classes: 150 A,H,PS Test Test rendering of the NASA NAIF spacecraft code list as on ontology.
NASA PDS4 Information Model Classes: 221 Properties: 114 PS Test Ontology version of the PDS4 Information Model
NASA PDS4 Information Model (Full dump) Classes: 1038 PS Test Test rendering of the NASA PDS4 Information Model as an ontology (full dump provided by the NASA team).
NASA PDS4 (from ESIP) Classes: 1149 PS Test PDS4 information mode rendered as on ontology by the ESIP group.
PDS Instrument Types Classes: 186 Individuals: 8 PS Test Test ontology of instrument types, from various IPDA sources.
SPASE Reference frames Classes: 47 H Develop Intermediate ontology extracted from the SPASE XML Schema. To be merged into IVOA RefFrame.
SPASE OWL Test Implementation Classes: 129 Individuals: 642 Properties: 273 M,H Develop Test rendering of the SPASE XML schema as an ontology.
Unified Astronomy Thesaurus (UAT) Concepts: 2373 A,H,PS Ready Used for keywords in astronomy papers
VAMDC Vocabulary Classes: 603 Properties: 554 M,A,H,PS Develop Test rendering of the VAMDC schema as an ontology.
Table 2: Subset of the list of semantic artefacts implemented in the astronomy OntoPortal instance prototyped during FAIR-IMPACT. The topics column letter codes are: A – Astronomy; H – Heliophysics; M – Metadata; PS – Planetary Sciences. The status can be: Test (for testing purposes); Develop (currently being developed); Ready (already in use).
4.3 Improvement of IVOA vocabularies
Importing the IVOA vocabularies into our astronomy OntoPortal instance revealed minor design issues in the IVOA vocabularies management and their associated RDF artefacts, as defined in Demleitner et al. (2023). We fixed the SKOS vocabularies, which were missing top level concepts. This prevented tools to correctly build the concept hierarchies out of the ontology. We also fixed the way the URI were referenced internally in the ontologies. Finally, we started to add more metadata to our vocabularies, such as, preferred namespace prefix and URIs, using the multiple properties harvested and promoted by MOD. We also initiated a mapping exercise between some of our properties with more commonly used ones. For instance, we have two properties used to handle deprecated terms and redirect to a replacement term: ivoasem:deprecated and ivoasem:useInstead, as part of our IVOA Semantics vocabulary (IVOASEM).
4.4 Strong heliophysics support
Thanks to the Transform to Open Science (TOPS) project, the NASA contributors to IHDEA started several projects aiming at improving semantic interoperability of the heliophysics community. A first example is the conversion of the SPASE XML Schema into an OWL ontology [^35] https://github.com/polyneme/topst-spase-rdf-tools. A second ongoing work is the mapping of the SPASE XML Schema to schema.org [^36] https://github.com/clnsmth/soso, in coordination with Science on Schema.org (Shepherd et al., 2024; Ringuette et al., 2025). When implemented and operational, these mappings will require an easy access to SAs, based search interfaces on ontology terms, so that term URIs are seamlessly inserted in the metadata. OntoPortal-Astro will provide a search interfaces for each SA, including a programmatic access.
5 Challenges and impact
As presented in the previous sections, the astronomy community is composed of several sub-communities (celestial astronomy, planetary sciences and heliophysics). Each community maintains and structures their SAs in heterogeneous forms. Since their drivers are also different, they have various maturity level concerning the management of SAs. They also store their ontologies in various repositories, using different technologies. However, the main goal of all the SAs is to enhance the FAIRness of research products (like, e.g., published datasets).
The OntoPortal technology has been identified to improve this situation, by bringing all SAs into a common catalogue with advanced capabilities. Nine impact factors have been identified by the FAIR-IMPACT team, which are directly linked to the adoption of the OntoPortal technology, and seven metrics are proposed to monitor the impact. The impact factors and metrics are listed in Figure 1.
Most of the proposed metrics are based on monitoring the evolution of SAs using OntoPortal-Astro: (a) number of expanded or emerging SA; (b) number of new SAs; (c) number of new mappings between SAs; and (d) the number of connection with other SA catalogues. The adoption of SAs by data alliances could also be assessed through OntoPortal-Astro, if the metadata include the list of projects the SAs are used in. The adoption of SAs in data repositories requires to browse the community repositories to assess the presence in the repository record metadata of URIs emanating from SAs referenced in OntoPortal-Astro. For the last two assessment methods, the developments done within the SKG-IF (Scientific Knowledge Graph Interoperability Framework) working group of the Research Data Alliance (RDA) will provide a way to access repositories and registries homogeneously, facilitating thus the monitoring of, e.g., dct:subjects URIs.
Figure 1: Impact factors and metrics for assessing the impact of the OntoPortal technology on the astronomy community. Figure adapted from Aubin et al. (2025).
6 Discussion
The development of the astronomy OntoPortal prototype has been truly interdisciplinary. It was enabled by the FAIR-IMPACT project, which has been funded European Open Science Cloud (EOSC) infrastructure programme. One of the goals of FAIR-IMPACT was to develop a multi-disciplinary federation of SA catalogues. The project selected the OntoPortal technology. Although the astronomy communities are mature communities, with well-established data ecosystems and solid bottom-up metadata governance, their use of semantic web technologies is not so much developed, compared to the practices in place in other communities like environmental or life sciences. The collaboration between the astronomy community on the one hand, and INRAE and the OntoPortal alliance on the other hand, brought astronomy the missing pieces to fast-forward to the modern age of web-semantics, with state-of-art technologies. The collaborations initiated during this project will be continued, with upcoming activities on OntoPortal-Astro and other projects of the EOSC programme.
As shown in Molinaro et al. (2020), in the frame of the ESCAPE project, the IVOA Registry has been connected to B2Find (a service provided by EUDAT, in the frame of EOSC): https://b2find.eudat.eu/organization/ivoa. This provides a new venue [^37] In the sense of the SKG-IF “venue” definition: A publishing “gateway” used by an Agent to make their Research products available to others. for finding IVOA services and collections. However, the current connection is not implementing any of the implicit IVOA semantics (like, e.g., the IVOA UAT keywords, which are available as pure text terms, and not URIs). The goal of the FAIR-IMPACT project was to expand FAIR solutions across EOSC, including semantic web technologies. Once OntoPortal-Astro is operational, further work with the IVOA and EUDAT/B2Find will have to be conducted to connect this new SA catalogue to the B2Find search portal.
The developments around the astronomy OntoPortal prototype have been regularly discussed within the IVOA Semantics WG. While the WG members are fully aware and positive about the open science challenges and the need for improving the astronomy SAs, the common ground is that the IVOA vocabularies are tailored for enabling interoperability within the IVOA. Hence, any development on the IVOA vocabularies shall be made backward compatible with the current uses and practices and shall not impose extra burden on the maintainers, tool developers and users of the vocabularies.
The IVOA vocabularies are setup so that they have as little external dependencies as possible. The main example is the UAT, which is developed independently from the IVOA. The Unified Astronomy Thesaurus (IVOA rendering) is a regularly updated copy of the UAT thesaurus. The IVOA version includes the relations to the original UAT terms, so that linked data applications can connect resources using the same terms. However, in the unfortunate event of discontinuation of the UAT reference thesaurus, the IVOA interfaces based on those terms will continue to work. The current design is thus trying to keep a strong infrastructure, as well as allowing linked and open data applications. For this, OntoPortal-Astro will serve as a long term solution in case a vocabulary cease to be supported (e.g., URIs are not resolvable anymore), as the technology include a URI resolution and negotiation service for each hosted SA.
Even with the impulse of OntoPortal-Astro, the semantic artefact management may not change in all communities immediately. For instance the IVOA has already a well established SA governance, and the improvement should only be limited to the improvement of the SA metadata. For planetary sciences and heliophysics, the first step will be to include OWL or SKOS versions of their SAs, and then include this scheme into their SA governance. Currently, the heliophysics community is studying the mapping of their registry metadata records to schema.org, which requires to produce SAs and accessible concept URIs. In the planetary science community, the IPDA is currently not ready for such a shift, but the planetary surfaces OGC group should be able to produce SAs related to coordinate systems and projections.
Improving the FAIRness of the astronomy vocabulary is also one of the key outcome of this study. We used FAIR assessment tools dedicated to ontologies and SAs, like FOOPS! [^38] https://foops.linkeddata.es/FAIR_validator.html (Garijo et al., 2021). Among all the assessment criteria, we can list the following aspects. The term URIs should be persistent, but the tools only consider a limited set of http domains as persistent. The metadata should include a list of contributors, for the ontology and the individual terms. It should also include provenance metadata. A versioning schema should also be in place, as discussed in Jonquet and Poveda-Villalón (2023). A preferred prefix should be advertised. Some criteria can easily be set up, and others (like the persistent identifier one) would require coordination with the FAIR assessment tool maintainers. On this aspect, the future OntoPortal-Astro will include O’FAIRe tool which provides a nice reporting on how to improve the FAIRness of each SAs.
This study is a first step towards improved management practices of SAs for astronomy. We focused here on vocabularies and ontologies, which are already in use by the astronomy communities. A further step is to ease and extend their usage for describing and annotating research objects. The astronomy community pilot developed within the OSTrails project will make use of the FAIR SAs developed using OntoPortal for connecting resources through Science Knowledge Graph interfaces. Finally, another framework identified for annotated research products with description of measurements is I-ADOPT. It is proposing a simple template with design patterns and is making use of existing ontologies. We participated to the I-ADOPT Variable Modelling Challenge in 2024, with an example variable “electron density in the solar wind”. From this simple test, the framework seems applicable, but, a lot concepts related to astronomy are missing. We will explore more examples with astronomy use cases and identify which ontologies and concepts should be implemented to cover the needs.
Keeping our astronomy ecosystem stable while opening up to linked data practices requires to build term mappings. Recently developed tools (see, e.g., MSCR – Metadata Schema and Crosswalk Registry, Kesäniemi et al., 2025) are promising resources for storing mappings, including more complex conditional schema mappings than the current capabilities of OntoPortal, as those developed during the SPASE XML Schema to Schema.org exercise. The mappings will then have to be stored, either in independent mapping registries (such as the MSCR or the native OntoPortal mapping repository) or integrated into our ontologies. The use of artificial intelligence and large language models will also be tested for aligning vocabularies and building mappings.
Since January 2025, the OSCARS cascading grant “Ontology Portal for Astronomy Linked-data” [^39] https://oscars-project.eu/projects/opal-ontology-portal-astronomy-linked-data (OPAL) started. Its goal is to build a real production service, OntoPortal-Astro for the astronomy community building on the preliminary results from FAIR-IMPACT. Through this project, we plan to gather, update and produce SAs for all the astronomy sub-communities. We will also explore how to better expose our SAs so that they can be reused. The OntoPortal-Astro team will include ontologists and knowledge designers, and will be supported by an Advisory and User Group composed of experts from the various astronomy communities. OntoPortal-Astro will join the Ontoportal Alliance, and the portal will federate with other OntoPortal instances, such as EarthPortal, which is serving ontologies we can relate our terms to (for the space weather community or the planetary surfaces). We also consider developing domain ontologies, which will allow us to annotate our research products in a more accurate manner.
7 Conclusion
OntoPortal-Astro may become the reference Semantic Artefact Catalogue for astronomy, it shall represent a significant step toward strengthening the semantic infrastructure in astronomy by providing a centralised service for FAIR astronomy SAs. By integrating and addressing the challenges from heliophysics to planetary sciences and across related astronomy disciplines, it will foster greater interoperability, data discoverability, and reuse. Through collaboration with the OntoPortal Alliance, this initiative will not only enhance the accessibility and management of SAs but also support interdisciplinary research across astronomy and beyond. Ultimately, OntoPortal-Astro will contribute to a more cohesive and FAIR-aligned semantic ecosystem, benefiting the broader scientific community.
Author Contributions
Baptiste Cecconi (IVOA, IPDA, IHDEA and FAIR-IMPACT): Writing - Original Draft, Conceptualization, Investigation, Methodology, Data Curation, Validation, Supervision. Laura Debisschop (FAIR-IMPACT): Writing - Review & Editing, Validation, Investigation, Data Curation. Sébastien Derrière (IVOA): Writing - Review & Editing, Conceptualization, Validation. Mireille Louys (IVOA): Writing - Review & Editing, Conceptualization, Validation. Clément Jonquet (OntoPortal, FAIR-IMPACT): Writing - Review & Editing, Project administration, Methodology, Software, Supervision. Carmen Corre (FAIR-IMPACT): Writing - Review & Editing, Methodology, Project administration. Nina Grau (FAIR-IMPACT): Writing - Review & Editing, Project administration, Methodology, Conceptualization, Data Curation.
Acknowledgments
This work has been supported by the FAIR-IMPACT project, which received funding from the European Commission’s Horizon Europe Research and Innovation programme under grant agreement number 101057344. The authors acknowledge the Observatoire de Paris-PSL for hosting the OntoPortal prototype used for this work and the OntoPortal Alliance for their support and technology. The authors also thank Markus Demleitner (former chair of the IVOA Semantics WG), as well as the IVOA, IHDEA and IPDA groups for their helpful discussions and feedback. The authors also acknowledge support from Sophie Aubin from INRAE.
Appendix A Semantic artefact namespaces
Table 1 lists the semantic artefact namespaces used throughout the paper.
SA Name Namespace URI Topics
IVOA Object Types Vocabulary objtyp https://ivoa.net/rdf/object-type/# A
EPNcore Target Classes tarcls https://voparis-ns.obspm.fr/rdf/epn/2.0/target-class/# P
Unified Astronomy Thesaurus uat http://astrothesaurus.org/uat/ A,P,H
IVOA UAT Vocabulary ivoauat %****␣semartastro.tex␣Line␣1225␣****https://ivoa.net/rdf/uat/# A,P,H
Simple Knowledge Organisation Systems (SKOS) skos http://www.w3.org/2004/02/skos/core# M
IVOA Semantics Vocabulary ivoasem https://ivoa.net/rdf/ivoasem/# M
Dublin Core Metadata Initiative (DCMI) Metadata Terms dct http://purl.org/dc/terms/ M
Table 1: List of namespaces used in this paper. The topics code letters are the same as in Table 2.